# On the multiplier spectrum of polynomials

### v1

@@ 2. Orbifolds and generalized Lattès maps
Let $A,B,\pi:\widehat{\mathbb{C}}\to\widehat{\mathbb{C}}$ be rational maps of degree at least $2$ with $A\circ\pi=\pi\circ B$. Then one of the following holds:
@@
    There exist non-constant rational maps $\psi,\pi_{0},B_{0}$ satisfying:
    • AA is a generalized Lattès map;
@@
    • $A:\mathcal{O}_{2}^{\pi_{0}}\to\mathcal{O}_{2}^{\pi_{0}}$ and $B_{0}:\mathcal{O}_{1}^{\pi_{0}}\to\mathcal{O}_{1}^{\pi_{0}}$ are minimal holomorphic maps between orbifolds;
@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
A particular class of rational maps, namely simple rational maps, plays a role in the proof of Theorem 1.3 in [JX25] (see [JX25, Theorem 3.3]). This notion was introduced by Pakovich [Pak25a]. A rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is simple if it has exactly 2d22d-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. By the Riemann–Hurwitz formula, every $f\in\mathbb{C}(z)$ of degree $d$ has exactly 2d22d-2 critical points counting with multiplicity. Hence, a simple rational map is a rational map with all critical points of multiplicity $1$ such that no two distinct critical points map to the same value.
For rational maps, simplicity is a natural notion since simple rational maps of degree $d\geq 2$ form a non-empty Zariski open subset of ${\rm Rat}_{d}(\mathbb{C})$. However, every polynomial $f\in\mathbb{C}[z]$ of degree $d\geq 2$ has a critical point of multiplicity d1d-1 at $\infty$. Thus, $f$ is simple if and only if $d=2$. The notion of simplicity is not useful for polynomials. To overcome this, we introduce a modified notion for polynomials:
@@
We show that most statements in [Pak25a] concerning simplicity for rational maps also hold for pre-simplicity for polynomials. In particular, we obtain a polynomial version (Theorem 3.23) of [JX25, Theorem 3.3], which yields a proof of Theorem 1.4 for $d\geq 4$. Most proofs in this section are minor modifications of those in [Pak25a]; we provide complete details for the reader’s convenience.
We first recall some notions about decomposing rational maps. A decomposition of a rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is a representation $f=f_{r}\circ\cdots\circ f_{1}$, where each $f_{i}$ has degree $\geq 2$. Two decompositions $f=f_{r}\circ\cdots\circ f_{1}$ and $f=g_{l}\circ\cdots\circ g_{1}$ are equivalent if either $l=r=1$ and $f_{1}=g_{1}$, or $l=r\geq 2$ and there exist Möbius transformations $\nu_{1},\dots,\nu_{r-1}\in{\rm PGL}_{2}(\mathbb{C})$ such that $f_{r}=g_{r}\circ\nu_{r-1}$, $f_{j}=\nu_{j}^{-1}\circ g_{j}\circ\nu_{j-1}$ for $1<j<r$, and $f_{1}=\nu_{1}^{-1}\circ g_{1}$. A rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is indecomposable if $f=f_{2}\circ f_{1}$ implies $\deg(f_{1})=1$ or $\deg(f_{2})=1$.
@@
A pre-simple polynomial $f$ of degree $d\geq 2$ is indecomposable (as a rational map) and satisfies ${\rm Mon}(f)\cong S_{d}$, where $S_{d}$ is the permutation group on $\{1,2,\dots,d\}$ and ${\rm Mon}(f)$ is the monodromy group of $f$.
@@
Let $f\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $d\geq 2$. Suppose $f$ is decomposable, so $f=f_{1}\circ f_{2}$ with $f_{1},f_{2}\in\mathbb{C}(z)$ of degrees $m_{1},m_{2}\geq 2$, respectively. Then $d=m_{1}m_{2}$. Since $\infty$ is a totally ramified fixed point of $f=f_{1}\circ f_{2}$, the point $a=f_{2}(\infty)$ satisfies $f_{1}^{-1}(\infty)=\{a\}$ and $f_{2}^{-1}(a)=\{\infty\}$. After composing with suitable Möbius transformations, we may assume $a=\infty$ and both $f_{1}$ and $f_{2}$ are polynomials. The pre-simplicity of $f$ implies that the number of critical values of $f$ in $\mathbb{C}$ is N(f)=(2d2)(d1)=d1N(f)=(2d-2)-(d-1)=d-1. On the other hand, from $f^{\prime}=(f_{1}^{\prime}\circ f_{2})\cdot f_{2}^{\prime}$, we have
@@
Equivalently, $(m_{1}-1)(m_{2}-1)\leq 0$, contradicting $m_{1},m_{2}\geq 2$. Hence, $f$ is indecomposable.
Since $f$ is indecomposable, its monodromy group ${\rm Mon}(f)$ is primitive. Fix a critical value $c$ of $f$ in $\mathbb{C}$; the corresponding permutation in ${\rm Mon}(f)$ is a transposition. By [Wie64, Theorem 13.3], ${\rm Mon}(f)$ must be the full symmetric group, so ${\rm Mon}(f)\cong S_{d}$. ∎
We now introduce some algebraic curves associated with polynomials, following [Pak25a, §2].
@@
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively. Define the algebraic curves $h_{F,H}$ and $h_{F}$ by
$h_{F,H}:F(x)-H(y)=0\quad\text{and}\quad h_{F}:\frac{F(x)-F(y)}{x-y}=0.$
The genera of these curves can be computed explicitly as follows; see [Fri74] or [Pak11].
Fact 3.4.
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively, and let $S=\{z_{1},\dots,z_{r}\}$ be the set of all critical values of $F$ or $H$ in $\mathbb{C}$. For $1\leq j\leq r$, let $(a_{j,1},\dots,a_{j,p_{j}})$ and $(b_{j,1},\dots,b_{j,q_{j}})$ denote the multiplicities of $F$ and $H$ at the points in $F^{-1}(z_{j})$ and $H^{-1}(z_{j})$, respectively. Then the genera of $h_{F,H}$ and $h_{F}$ satisfy
$2-2g(h_{H,F})=\gcd(m,n)-(r-1)mn+\sum_{j=1}^{r}\sum_{s_{1}=1}^{p_{j}}\sum_{s_{2}=1}^{q_{j}}\gcd(a_{j,s_{1}},b_{j,s_{2}})$
and
$4-2g(h_{F})=m-(r-1)m^{2}+\sum_{j=1}^{r}\sum_{s_{1}=1}^{p_{j}}\sum_{s_{2}=1}^{p_{j}}\gcd(a_{j,s_{1}},a_{j,s_{2}}).$
The following two theorems are analogues of [Pak25a, Theorems 2.3 and 2.4].
@@
Let $m\geq 4$ and $n\geq 2$ be integers, with $n\neq 2,4$ when $m=4$. Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m$ and $H\in\mathbb{C}[z]$ a polynomial of degree $n$. Assume the curve $h_{F,H}$ is irreducible. Then $g(h_{F,H})>0$. In particular, there exist no non-constant rational functions $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$ in $\mathbb{C}(z)$.
@@
We use the notation from Fact 3.4. Let ${\rm critV}(F)$ and ${\rm critV}(H)$ denote the sets of critical values of $F$ and $H$ in $\mathbb{C}$, respectively.
For $i\in\{1,\dots,r\}$ with $z_{i}\notin{\rm critV}(F)$, we have $p_{i}=m$ and $(a_{i,1},\dots,a_{i,m})=(1,\dots,1)$, so
@@
For $i\in\{1,\dots,r\}$ with $z_{i}\in{\rm critV}(F)$, since $F$ is pre-simple, we have $\deg_{z}(F)\in\{1,2\}$ for every $z\in\mathbb{C}$. Then $p_{i}=m-1$ and we may assume $(a_{i,1},\cdots,a_{i,m-2},a_{i,m-1})=(1,\dots,1,2)$. Let $l_{i}$ be the number of indices $j_{2}\in\{1,\dots q_{i}\}$ such that $b_{i,j_{2}}$ is even. Then
@@
$\sum_{i=1}^{r}\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}})=\sum_{i=1}^{r}mq_{i}+\Sigma=(r-1)mn+m+\Sigma,$
where $\Sigma:=\sum_{1\leq i\leq r,\,z_{i}\in{\rm critV}(F)}(l_{i}-q_{i})$. By Fact 3.4, the genus is
$g(h_{H,F})=\frac{1}{2}\cdot(2-m-\gcd(m,n)-\Sigma).$
Hence, $g(h_{H,F})=0$ if and only if $\Sigma=m+\gcd(m,n)-2$. Note that $\Sigma$ equals the number of $z\in\mathbb{C}$ such that $H(z)\in{\rm critV}(F)$ and $\deg_{z}(H)$ is odd.
For any finite subset $A\subset\mathbb{C}$, we have $n-1\geq\sum_{z\in H^{-1}(A)}(\deg_{z}(H)-1)$; hence $H^{-1}(A)\geq n(\#A)-(n-1)=n(-1+\#A)+1$, with equality if and only if ${\rm critV}(H)\subseteq S$. In particular, $H^{-1}({\rm critV}(F))\geq n(m-2)+1$, with equality if and only if ${\rm critV}(H)\subseteq{\rm critV}(F)$.
Suppose $g(h_{F,H})=0$. Then $\Sigma=m+\gcd(m,n)-2$, which implies
$\#H^{-1}({\rm critV}(F))$
@@
$=$ $\frac{1}{2}((m-1)(n+1)+\gcd(m,n)-1).$
Thus, $(m-1)(n+1)+\gcd(m,n)-1\geq 2n(m-2)+2$, or equivalently, $mn+4\leq 3n+m+\gcd(m,n)$. Since $\gcd(m,n)\leq n$, we get $mn+4\leq 4n+m$, so $(m-4)(n-1)\leq 0$. Given $n\geq 2$ and $m\geq 4$, we must have $m=4$. Then $n\neq 2,4$ by assumption. Now $mn+4\leq 3n+m+\gcd(m,n)$ becomes $n\leq\gcd(4,n)$, contradicting $n\neq 2,4$. Therefore, $g(h_{H,G})>0$.
Now suppose there exist $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$. Let $C$ be the projective curve in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by $h_{F,H}$, which is irreducible by assumption. Then $\phi=(X,Y):\Delta\dashrightarrow C$ is a non-constant rational map, where $\Delta$ is the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$. Since $\Delta\cong\mathbb{P}^{1}$ is non-singular and $C$ is projective, $\phi:\Delta\to C$ is a morphism. But $g(\Delta)=0$ and $g(C)>0$, so by the Riemann–Hurwitz formula, no non-constant morphism from $\Delta$ to $C$ exists, hence we get a contradiction. ∎
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 4$. Then the curve $h_{F}$ is irreducible and $g(h_{F})>0$. In particular, there exist no non-constant rational functions $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=F\circ Y$ in $\mathbb{C}(z)$, except when $X=Y$.
@@
By Proposition 3.2, ${\rm Mon}(F)\cong S_{m}$. In particular, ${\rm Mon}(F)$ is doubly transitive, so the curve $h_{F}$ is irreducible by [Pak11, Corollary 2.3]. By Fact 3.4, since $f$ is pre-simple, we have
$4-2g(h_{F})=m-(m-1-1)m^{2}+(m-1)((m-2)(m-1)+(m-2+2))=-m^{2}+5m-2,$
i.e., $g(h_{F})=(m-2)(m-3)/2$. For $m\geq 4$, we have $g(h_{F})>0$.
Suppose there exist $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ with $X\neq Y$ such that $F\circ X=F\circ Y$. Let $C$ be the irreducible projective curve in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ given by $h_{F}$. Since $X\neq Y$, the map $\phi=(X,Y):\Delta\dashrightarrow C$ is a non-constant rational map, where $\Delta$ is the diagonal in $\mathbb{P}^{1}\times\mathbb{P}^{1}$. As before, $\phi:\Delta\to C$ is a morphism. But $g(\Delta)=0$ and $g(C)>0$, leading to a contradiction by the Riemann–Hurwitz formula. ∎
Self-compositions of a pre-simple polynomial behave well under decomposition.
Theorem 3.7.
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Then for every $l\in\mathbb{Z}_{>0}$, every decomposition of $F^{\circ l}$ into a composition of indecomposable rational maps (of degree $\geq 2$) is equivalent to the decomposition $F^{\circ l}=F\circ\cdots\circ F$.

### v2

@@ 2. Orbifolds and generalized Lattès maps
Let $A,B,\pi\in\mathbb{C}(z)$ be rational maps of degree at least $2$ such that $A\circ\pi=\pi\circ B$. Then one of the following holds:
@@
    AA is a generalized Lattès map and there exist $\psi,\pi_{0},B_{0}\in\mathbb{C}(z)\setminus\mathbb{C}$ satisfying:
@@
    • Both $A:\mathcal{O}_{2}^{\pi_{0}}\to\mathcal{O}_{2}^{\pi_{0}}$ and $B_{0}:\mathcal{O}_{1}^{\pi_{0}}\to\mathcal{O}_{1}^{\pi_{0}}$ are minimal holomorphic maps between orbifolds;
@@ 3.1. Pre-simple polynomials and the case $d\geq 4$
A particular class of rational maps, namely simple rational maps, plays a key role in the proof of Theorem 1.3 in [JX25] (see [JX25, Theorem 3.3]). This notion was introduced by Pakovich [Pak25a]. A rational map $f\in\mathbb{C}(z)$ of degree $d\geq 2$ is simple if it has exactly 2d22d-2 distinct critical values in $\mathbb{P}^{1}(\mathbb{C})$. By the Riemann–Hurwitz formula, every $f\in\mathbb{C}(z)$ of degree $d$ has exactly 2d22d-2 critical points counted with multiplicity. Hence, a simple rational map is a rational map with all critical points of multiplicity $1$ such that no two distinct critical points map to the same value.
For rational maps, simplicity is a natural notion since simple rational maps of degree $d\geq 2$ form a Zariski open and dense subset of $\mathrm{Rat}_{d}(\mathbb{C})$. However, every polynomial $f\in\mathbb{C}[z]$ of degree $d\geq 2$ has a critical point of multiplicity d1d-1 at $\infty$. Thus, $f$ is simple if and only if $d=2$. The notion of simplicity is thus ill-suited for polynomials. To overcome this, we introduce a modified notion for polynomials:
@@
Note that d1d-1 is the maximum possible number of distinct finite critical values for a polynomial of degree $d$. We show that most statements in [Pak25a] concerning simple rational maps admit analogues for pre-simple polynomials. In particular, we obtain a polynomial version (Theorem 3.29) of [JX25, Theorem 3.3], which yields a proof of Theorem 1.4 for $d\geq 4$. Most proofs in this subsection are minor modifications of those in [Pak25a]; we provide complete details for the reader’s convenience.
We first recall some notions about decomposing rational maps. Let $f\in\mathbb{C}(z)$ be a rational map of degree $d\geq 2$. We say that $f$ is indecomposable if $f=f_{1}\circ f_{2}$ in $\mathbb{C}(z)$ implies $\deg(f_{1})=1$ or $\deg(f_{2})=1$. A decomposition of $f$ is a representation $f=f_{1}\circ\cdots\circ f_{r}$, where each $f_{i}\in\mathbb{C}(z)$ has degree $\geq 2$. Two decompositions $f=f_{1}\circ\cdots\circ f_{r}$ and $f=g_{1}\circ\cdots\circ g_{l}$ of $f$ are called equivalent if either $l=r=1$ and $f_{1}=g_{1}$, or $l=r\geq 2$ and there exist $\nu_{1},\dots,\nu_{r-1}\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $f_{1}=g_{1}\circ\nu_{1}$, $f_{j}=\nu_{j-1}^{-1}\circ g_{j}\circ\nu_{j}$ for $1<j<r$, and $f_{r}=\nu_{r-1}^{-1}\circ g_{r}$. A decomposition $f=f_{1}\circ\cdots\circ f_{r}$ of $f$ is called complete if $f_{j}$ is indecomposable for every $1\leq j\leq r$. It is easy to see that any rational map of degree $d\geq 2$ admits a complete decomposition by an induction on $d$.
@@
A pre-simple polynomial $f$ of degree $d\geq 2$ is indecomposable (as a rational map) and satisfies $\mathrm{Mon}(f)\cong S_{d}$, where $S_{d}$ is the permutation group on $\{1,2,\dots,d\}$ and $\mathrm{Mon}(f)$ is the monodromy group of $f$.
@@
Let $f\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $d\geq 2$. Suppose $f$ is decomposable, so $f=f_{1}\circ f_{2}$ with $f_{1},f_{2}\in\mathbb{C}(z)$ of degrees $m_{1},m_{2}\geq 2$, respectively. Then $d=m_{1}m_{2}$. Since $\infty$ is a totally invariant fixed point of $f=f_{1}\circ f_{2}$, the point $a=f_{2}(\infty)$ satisfies $f_{1}^{-1}(\infty)=\{a\}$ and $f_{2}^{-1}(a)=\{\infty\}$. After composing with a suitable Möbius transformation, we may assume $a=\infty$. Then both $f_{1}$ and $f_{2}$ are polynomials. The pre-simplicity of $f$ implies that the number $N(f)$ of critical values of $f$ in $\mathbb{C}$ equals (2d2)(d1)=d1(2d-2)-(d-1)=d-1. On the other hand, from $f^{\prime}=(f_{1}^{\prime}\circ f_{2})\cdot f_{2}^{\prime}$ we see that
@@
Equivalently, $(m_{1}-1)(m_{2}-1)\leq 0$, contradicting the assumption that $m_{1},m_{2}\geq 2$. Hence, $f$ is indecomposable.
Since $f$ is indecomposable, its monodromy group $\mathrm{Mon}(f)$ is primitive. Fix a critical value $c$ of $f$ in $\mathbb{C}$. Since $f$ is pre-simple, the fiber $f^{-1}(c)$ contains exactly one critical point with local degree $2$ and d2d-2 unramified points; hence, the corresponding permutation in $\mathrm{Mon}(f)$ is a transposition. By [Wie64, Theorem 13.3], $\mathrm{Mon}(f)$ must be the full symmetric group, so $\mathrm{Mon}(f)\cong S_{d}$. ∎
We now introduce algebraic curves associated with polynomial pairs, following [Pak25a, § 2].
@@
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively. Let $h_{F,H}$ be the algebraic curve (in $\mathbb{A}^{2}$) defined by F(x)H(y)=0F(x)-H(y)=0.
The genus of the curve $h_{F,H}$ can be computed explicitly as follows (see [Fri74] or [Pak11]):
Proposition 3.4.
Let $F(z),H(z)\in\mathbb{C}[z]$ be polynomials of degrees $m,n\geq 1$, respectively, and let $S=\{z_{1},\dots,z_{r}\}$ be the union of the sets of critical values of $F$ and $H$ in $\mathbb{C}$. For $1\leq i\leq r$, let $(a_{i,1},\dots,a_{i,p_{i}})$ and $(b_{i,1},\dots,b_{i,q_{i}})$ denote the multiplicities of $F$ and $H$ at the points in $F^{-1}(z_{i})$ and $H^{-1}(z_{i})$, respectively. Then the genus of $h_{F,H}$ is determined by
$2-2g(h_{F,H})=\gcd(m,n)-(r-1)mn+\sum_{i=1}^{r}\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}}).$
The following theorem is an analogue of [Pak25a, Theorem 2.3].
@@
Let $m\geq 4$ and $n\geq 2$ be integers such that $(m,n)$ is not $(4,2)$ or $(4,4)$. Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m$ and $H\in\mathbb{C}[z]$ a polynomial of degree $n$. Assume the curve $h_{F,H}$ is irreducible. Then $g(h_{F,H})>0$. In particular, there exist no non-constant rational maps $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$ in $\mathbb{C}(z)$.
@@
We use the notation from Proposition 3.4. Let $\mathrm{CritV}(F)$ and $\mathrm{CritV}(H)$ denote the sets of critical values of $F$ and $H$ in $\mathbb{C}$, respectively.
For $i\in\{1,\dots,r\}$ such that $z_{i}\notin\mathrm{CritV}(F)$, we have $p_{i}=m$ and $(a_{i,1},\dots,a_{i,m})=(1,\dots,1)$, so
@@
For $i\in\{1,\dots,r\}$ such that $z_{i}\in\mathrm{CritV}(F)$, since $F$ is pre-simple, we have $p_{i}=m-1$ and $\deg_{z}(F)\in\{1,2\}$ for every $z\in\mathbb{C}$. We may assume
$(a_{i,1},\dots,a_{i,m-2},a_{i,m-1})=(1,\dots,1,2).$
Let $l_{i}$ be the number of indices $j_{2}\in\{1,\dots,q_{i}\}$ such that $b_{i,j_{2}}$ is even. Then
@@
$\sum_{i=1}^{r}\sum_{j_{1}=1}^{p_{i}}\sum_{j_{2}=1}^{q_{i}}\gcd(a_{i,j_{1}},b_{i,j_{2}})=\sum_{i=1}^{r}mq_{i}+A=(r-1)mn+m+A,$
where $A:=\sum_{1\leq i\leq r,\,z_{i}\in\mathrm{CritV}(F)}(l_{i}-q_{i})$. By Proposition 3.4, the genus is
$g(h_{F,H})=\frac{1}{2}\cdot(2-m-\gcd(m,n)-A).$
Hence, $g(h_{F,H})=0$ if and only if $A=m+\gcd(m,n)-2$. Note that A-A equals the number of $z\in\mathbb{C}$ such that $H(z)\in\mathrm{CritV}(F)$ and $\deg_{z}(H)$ is odd.
For any finite subset $S\subset\mathbb{C}$, we have
$n-1\geq\sum_{z\in H^{-1}(S)}(\deg_{z}(H)-1)=n(\#S)-\#H^{-1}(S);$
hence $H^{-1}(S)\geq n(\#S)-(n-1)=n(-1+\#S)+1$, with equality if and only if $\mathrm{CritV}(H)\subseteq S$. In particular, $H^{-1}(\mathrm{CritV}(F))\geq n(m-2)+1$, with equality if and only if $\mathrm{CritV}(H)\subseteq\mathrm{CritV}(F)$.
Suppose $g(h_{F,H})=0$. Then $A=m+\gcd(m,n)-2$, which implies
$\#H^{-1}(\mathrm{CritV}(F))$
@@
$=$ $\frac{1}{2}\left((m-1)(n+1)+\gcd(m,n)-1\right).$
Thus,
$(m-1)(n+1)+\gcd(m,n)-1\geq 2\#H^{-1}(\mathrm{CritV}(F))\geq 2n(m-2)+2,$
or equivalently,
$mn+4\leq 3n+m+\gcd(m,n).$
Then $mn+4\leq 3n+m+n$, so $(m-4)(n-1)\leq 0$. As $n\geq 2$ and $m\geq 4$, we must have $m=4$. Then $n\neq 2,4$ by assumption. Now $mn+4\leq 3n+m+\gcd(m,n)$ becomes $n\leq\gcd(4,n)$, contradicting the assumption that $n\neq 2,4$. Therefore, $g(h_{F,H})>0$.
Now suppose there exist $X(z),Y(z)\in\mathbb{C}(z)\setminus\mathbb{C}$ such that $F\circ X=H\circ Y$. Let $C$ be the Zariski closure of $h_{F,H}$ in $\mathbb{P}^{1}\times\mathbb{P}^{1}$ , which is an irreducible projective curve. Since genus is a birational invariant, $g(C)=g(h_{F,H})>0$. The morphism $(X,Y):\mathbb{P}^{1}\to\mathbb{P}^{1}\times\mathbb{P}^{1},z\mapsto(X(z),Y(z))$ has image contained in $C$. So it induces a non-constant morphism $\phi:\mathbb{P}^{1}\to C$ between irreducible projective curves. However, as $g(\mathbb{P}^{1})=0$ and $g(C)>0$, no non-constant morphism $\mathbb{P}^{1}\to C$ exists by the Riemann–Hurwitz formula, which is a contradiction. ∎
For a non-constant rational map $f\in\mathbb{C}(z)$, define the groups
$\Sigma(f):=\{\sigma\in\mathrm{PGL}_{2}(\mathbb{C})\colon f\circ\sigma=f\}\quad\text{and}\quad\Sigma_{\infty}(f):=\bigcup_{k\geq 1}\Sigma(f^{\circ k}).$
Clearly, $\Sigma(f)\subseteq\Sigma_{\infty}(f)$, and this inclusion is preserved under conjugacy in the sense that for $\tau\in\mathrm{PGL}_{2}(\mathbb{C})$,
(3.1) $\Sigma(\tau\circ f\circ\tau^{-1})=\tau\circ\Sigma(f)\circ\tau^{-1}\subseteq\tau\circ\Sigma_{\infty}(f)\circ\tau^{-1}=\Sigma_{\infty}(\tau\circ f\circ\tau^{-1}).$
For a polynomial $f\in\mathbb{C}[z]\setminus\mathbb{C}$, the totally invariant fixed point $\infty$ of $f^{\circ k}$ ($k\geq 1$) implies that
(3.2) $\Sigma_{\infty}(f)\subseteq\mathrm{Aff}(\mathbb{C}).$
Lemma 3.6.
Let $f\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Then $\Sigma(f)=1$.
Proof.
Since $f$ is pre-simple of degree $m\geq 3$, $f$ is not conjugate to $z^{m}$. After conjugacy, we may assume that $f$ is of the form
$f(z)=z^{m}+a_{m-2}z^{m-2}+\cdots+a_{0},$
where $a_{m-2},\dots,a_{0}\in\mathbb{C}$. Let $\tau\in\Sigma(f)$ and write $\tau(z)=\lambda z+c$ with $\lambda\in\mathbb{C}^{\times}$ and $c\in\mathbb{C}$ by (3.2). Comparing the coefficients of $z^{m}$ and $z^{m-1}$, we see that $c=0$ and $\lambda^{m}=1$. Assume that $\lambda\neq 1$. Let $l\geq 2$ be the order of the root of unity $\lambda$ in $\mathbb{C}^{\times}$. Then $l\mid m$. As $f(\lambda z)=f(z)$, we have $f(z)=g(z^{l})$ for a unique polynomial $g\in\mathbb{C}[z]$ of degree $m/l$. Since $f$ is pre-simple, its derivative $f^{\prime}(z)=lz^{l-1}g^{\prime}(z^{l})$ has no multiple roots. In particular, 0 is not a multiple root of $f^{\prime}$; hence $l=2$ (otherwise $z^{2}\mid f^{\prime}(z)$). Note that $f^{\prime}(z)=2zg^{\prime}(z^{2})\in\mathbb{C}[z]$ is a polynomial of degree $m-1\geq 2$ that is an odd function on $\mathbb{C}$ without multiple roots. Hence there exists $\omega_{0}\in\mathbb{C}^{\times}$ such that $f^{\prime}(\omega_{0})=f^{\prime}(-\omega_{0})=0$. However, the two distinct critical points $\omega_{0}\neq-\omega_{0}$ of $f$ are mapped by $f$ to the same value $g(\omega_{0}^{2})$, contradicting the assumption that $f$ is pre-simple. This completes the proof. ∎
The following theorem (see also Remark 3.8) is an analogue of [Pak25a, Theorem 2.4].
@@
Let $F\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Then the equality $F\circ X=F\circ Y$ in $\mathbb{C}[z]$ with $X,Y\in\mathbb{C}[z]\setminus\mathbb{C}$ implies $X=Y$.
@@
Let $X,Y\in\mathbb{C}[z]$ be non-constant polynomials such that $F\circ X=F\circ Y$. Set $n=\deg(X)=\deg(Y)\in\mathbb{Z}_{>0}$.
By [FG22, Theorem 3.54], there exist polynomials
$\tilde{F},\mu_{1},\mu_{2},G,\sigma_{1},\sigma_{2}\in\mathbb{C}[z]\setminus\mathbb{C}$
with
$\deg(\tilde{F})=m\quad\text{and}\quad\deg(G)=n$
such that
$F=\tilde{F}\circ\mu_{1}=\tilde{F}\circ\mu_{2},\quad\sigma_{1}\circ G=X,\quad\sigma_{2}\circ G=Y,\quad\text{and}\quad\mu_{1}\circ\sigma_{1}=\mu_{2}\circ\sigma_{2}.$
Note that $\mu_{1}$, $\mu_{2}$, $\sigma_{1}$, and $\sigma_{2}$ are polynomials of degree one. Then $F\circ\mu_{1}^{-1}\circ\mu_{2}=F$ and we deduce that $\mu_{1}=\mu_{2}$ by Lemma 3.6. Thus $\sigma_{1}=\mu_{1}^{-1}\circ\mu_{2}\circ\sigma_{2}=\sigma_{2}$. So $X=\sigma_{1}\circ G=\sigma_{2}\circ G=Y$. The proof is complete. ∎
Remark 3.8.
The presented proof of Theorem 3.7 is based on Ritt’s theory of polynomial decompositions.
If the pre-simple polynomial $F\in\mathbb{C}[z]$ has degree $m\geq 4$, then we can show that the algebraic curve $h_{F}$ in $\mathbb{A}^{2}$ given by
$\frac{F(x)-F(y)}{x-y}=0$
is irreducible and has genus $g(h_{F})>0$, using the argument in the proof of [Pak25a, Theorem 2.4]. Hence for pre-simple $F\in\mathbb{C}[z]$ of degree $m\geq 4$, we conclude that the equality $F\circ X=F\circ Y$ in $\mathbb{C}(z)$ with $X,Y\in\mathbb{C}(z)\setminus\mathbb{C}$ implies $X=Y$.
Using Theorem 3.7, we obtain the following generalization of Lemma 3.9:
Corollary 3.9.
Let $f\in\mathbb{C}[z]$ be a pre-simple polynomial of degree $m\geq 3$. Then $\Sigma(f)=\Sigma_{\infty}(f)=1$.